Calculate the maximum concentration (in M) of silver ions (Ag+) in a solution that contains 0.00125 M of CO32-. The Ksp of Ag2CO3 is 8.46×10-12.
Answer #1

44 ng co - 2 Ag+ con C11 Ksp - (Ag)] = kse/ (607) 8.464 10 12 0.00125 (Ag) = 676871012 (Agt] - 1676841522 (Agt) = 8.22+105